answer all questions

2.08 points kr mathoda (Da not round intermediate calculations. Round your final answar to 2 dacimal placas.)
Answer #1

Solution:

Amount invested = $8,860

Required amount after 18 years = $57,906

Let rate of return earned = i%

Now

P * (1+i)^n = $57,906

$8,860 (1+i)^18 = $57,906

(1+i)^18 = 6.5356659

1+i = (6.5356659)^1/18

i = 10.99%